OpenStudy (anonymous):

Independence Day question!! X3 Who knows how fireworks work? What makes them change color? Why? Have Fun!

OpenStudy (anonymous):

By the way, no googleing!

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

it contains different elements, which on burning produces characteristic colours.

OpenStudy (anonymous):

Well, I was sort of thinking something more in depth. I personally know the answer, but I thought it would be fun to see what people would answer. X3

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

And I also noticed this is chemistry and not Biology. Oh well... X3

OpenStudy (abhisar):

You probably didn't notice, but you posted your \(\large\sf\color{blue}{\underline{\href{ /study#/groups/Chemistry }{Chemistry }}}\) question in the Biology section. Next time when you ask a question, check if you're in the right section :)

OpenStudy (anonymous):

Yup. I know. I already said that... Hahahahahahahahahahha

OpenStudy (abhisar):

It's spamming according to \(\large\sf\color{red}{\underline{\href{ /code-of-conduct}{Code~of~conduct}}}\)

OpenStudy (anonymous):

Seriously. I was just trying to do something nice...

OpenStudy (anonymous):

"Am I not good enough?!?!?!?!" -Lady bug from Bee and Puppycat

OpenStudy (anonymous):

@emmigrace222

OpenStudy (abhisar):

repost it in \(\large\sf\color{blue}{\underline{\href{ /study#/groups/Chemistry }{Chemistry }}}\) section \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (anonymous):

Wait is there an option for that?

OpenStudy (abhisar):

Close it and post in chem ! :)

OpenStudy (anonymous):

Aww... okay....

OpenStudy (abhisar):

take it in a positive way :)

OpenStudy (emmigrace222):

hmmmm